The maximum current power in the external battery circuit at a current of 0.2A is 4W.find the emf and the resistance of the battery
1
Expert's answer
2020-11-17T07:10:26-0500

Maximum power given "(P) =4W"


Current at Maximum power "(I) =0.2A"


The load resistance "(R) =\frac{P}{I^2}=\frac{4}{0.2^2}=100\varOmega"


According to maximum power transfer theorm, during maximum power transfer load resistance =source resistance


i.e Source resistance"(Rs) =100\varOmega"


Emf of battery"(E) =I(R+Rs) =0.2\times200=40V"